# What Do You See? Enhancing Zero-Shot Image Classification with Multimodal Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have been effectively used for many computer vision tasks, including image classification. In this paper, we present a simple yet effective approach for zero-shot image classification using multimodal LLMs. By employing multimodal LLMs, we generate comprehensive textual representations from input images. These textual representations are then utilized to generate fixed-dimensional features in a cross-modal embedding space. Subsequently, these features are fused together to perform zero-shot classification using a linear classifier. Our method does not require prompt engineering for each dataset; instead, we use a single, straightforward, set of prompts across all datasets. We evaluated our method on several datasets, and our results demonstrate its remarkable effectiveness, surpassing benchmark accuracy on multiple datasets. On average over ten benchmarks, our method achieved an accuracy gain of 4.1 percentage points, with an increase of 6.8 percentage points on the ImageNet dataset, compared to prior methods. Our findings highlight the potential of multimodal LLMs to enhance computer vision tasks such as zero-shot image classification, offering a significant improvement over traditional methods.
1 Introduction
@@
Zero-shot image classification aims to categorize images into classes unseen during training, presenting a significant challenge in computer vision. Recent approaches leverage the power of large language models (LLMs) like GPT-4 (Brown et al. (2020)) to generate prompts for target classes, often in conjunction with vision-language models such as CLIP (Radford et al. (2021)) to embed images and text in a common space. Open-vocabulary models like CLIP (Radford et al. (2021)) and VirTex (Desai & Johnson (2021)) have shown promise in this area due to their ability to generalize to unseen classes. These models learn to match images with captions from vast amounts of image-text data, allowing for dynamic classification without retraining. Early works like DeViSE (Frome et al. (2013)) pioneered the concept of a joint embedding space for images and text, enabling generalization to unseen classes. Approaches like CLIP (Radford et al. (2021)), based on contrastive learning, and ALIGN (Jia et al. (2021)), employing a two-stage framework, further refined the alignment of image and text representations. More recent approaches for zero-shot image classification (e.g., Pratt et al. (2023); Menon & Vondrick (2023)) have utilized LLMs to generate prompts (i.e., captions or descriptions) for the target classes to further improve the classification accuracy.
@@
However, relying solely on visual features of the input images during inference can limit accuracy, as these features may not be sufficient to fully capture the nuances present in textual descriptions. To address this, we propose a novel method that leverages the capabilities of multimodal LLMs to generate rich textual representations of the input images. Multimodal LLMs, such as GPT-4 (Brown et al. (2020)) and Gemini (Gemini Team Google (2023)), have demonstrated remarkable abilities in various tasks. They are capable of processing and integrating information from various sources like text, images, and audio. This allows them to perform tasks that were previously impossible, such as generating detailed image descriptions, answering complex visual questions, and even creating realistic images from text. Inspired by these advancements, we utilize a straightforward set of prompts to generate detailed textual descriptions of the input images, eliminating the need for complex prompt engineering seen in previous works (e.g., Radford et al. (2021); Guo et al. (2023)). These textual representations are then fused with visual features to perform zero-shot classification. See Figure 1.
Our method offers several key advantages: it significantly improves classification accuracy by incorporating richer textual information extracted directly from the input images; it employs a simple and universal set of prompts, eliminating the need for dataset-specific prompt engineering; and it outperforms existing methods on a variety of benchmark datasets. By employing multimodal LLMs and a straightforward set of prompts, our method outperforms previous zero-shot image classification methods. Specifically, our method achieves an average accuracy gain of 4.1 percentage points over ten image classification benchmark datasets and an accuracy increase of 6.8% on the ImageNet dataset (Deng et al. (2009)).
@@
In the following sections, we detail our proposed approach for zero-shot image classification using LLMs (Section 2), present experimental results across ten benchmark datasets (Section 3), analyze the computational resources used (Section 4), discuss limitations (Section 5), and conclude with future directions (Section 6).
@@ 2 Method
Given an input image, X, containing object(s) belonging to a single class label from a finite set of class labels, , our objective is to classify X without any dataset-specific training process for image classification. The overview of our method is illustrated in Figure 2. As shown in Figure 2, our approach relies on a cross-modal embedding encoder models (image encoder, , and text encoder, ), trained to learn joint representations of images and text, as demonstrated in prior works, such as Radford et al. (2021); Frome et al. (2013); Jia et al. (2021). Additionally, we utilize a multimodal LLM, , which is pre-trained on a large corpus of multimodal data. This model, , is designed to generate responses that align with both textual and visual inputs, effectively integrating information from both modalities for enhanced predictions (e.g., Gemini Team Google (2023)).
@@ 2.1 Class label features
Our zero-shot classifier utilizes class label features of the target dataset. We first encode the set of class labels, , in some embedding space by the cross-modal encoder models, , and , such that refers to the set of encoded class label features, where is a normalized -D class label feature of the class label and is the dimensionality of the embedded features.
@@
Such an encoded feature matrix can be generated using one of three options: (1) directly from the textual class labels, (2) using a human-designed template – e.g., ‘‘A photo of {class_label}” (Radford et al. (2021); Guo et al. (2023)), where {class_label} refers to the textual label of one of our classes , or (3) LLM-generated class description(s) (Pratt et al. (2023)). In option (3), the LLM-generated class description(s) are then converted to embedded features, followed by fusion (e.g., averaging) to generate a single embedded feature for each class label in the dataset. Optionally, all features from the three options can be fused together (e.g., averaged) for increased robustness.
@@ 2.2 Cross-modal input features
To predict the final class, we first encode the input image by the cross-modal image encoder model, , to generate the image feature as described below:
where performs vector normalization. Traditionally, the image feature serves as the sole input to prior zero-shot image classifiers (Pratt et al. (2023); Radford et al. (2021); Guo et al. (2023)). Our method enhances this input by incorporating the LLM, , which generates additional textual-based inputs for our zero-shot image classifier (see Figure 3). To achieve this, we employ an engineered prompt that instructs the LLM to describe the input image, X, and perform initial image classification using the textual names of the class label set . Denoting and as our prompts for image description and initial image classification, respectively, we generate two additional embedded features alongside by:
where and refer to the image description feature and initial class prediction feature, respectively. Notably, in our method, unlike prior methods (e.g., Pratt et al. (2023); Radford et al. (2021); Guo et al. (2023)), such prompts and do not require dataset-specific engineering for each dataset. Instead, we employ fixed prompts: the first prompt instructs the LLM to provide a generic image description, while the second prompt includes the textual class labels of the target dataset. Further details regarding our prompts are provided in the supplemental materials (Appendix A).
After generating the three input features (image feature, description feature, and initial prediction feature), we fuse them to generate our final query feature, . One can interpret this fusion as an ensemble of different candidate features to generate a more precise query feature (see Figure 4). We adopted a simple fusion, where the final query feature, , is generated by:
@@ 2.3 Class label prediction
where represents the vector transpose operation to transform into a row vector of shape , and contains the similarity scores of the generated query feature to the class label features in the target dataset. The index of the final predicted class is then computed as that corresponds to the maximum similarity score.
@@ 3 Experiments
In our experiments, we employed Gemini Pro (Gemini Team Google (2023)) as our multimodal LLM, , for generating image descriptions and initial predictions. We utilized CLIP (ViT-L/14) (Radford et al. (2021)) as our cross-modal embedding encoder models, and , to encode input testing images, image descriptions, and initial predictions generated by Gemini Pro.
We explored four different versions of class label features, M, in Equation 1. Specifically, we used CLIP to encode the following representations of class labels: 1) class label names, 2) the text template ‘‘A photo of {class_label}’’, where {class_label} denotes each class label in each dataset, and 3) class descriptions, similar to those generated by CuPL (Pratt et al. (2023)), and 4) a combination of the aforementioned three options—akin to the fusion of our input features, we combined the three encoded features of each class and computed their average feature. The class descriptions were produced by prompting Gemini Pro to describe each class label in the dataset 50 times, resulting in 50 different class descriptions for each class. Subsequently, we utilized CLIP to encode the 50 class descriptions for each class and compute the average encoded feature vector to represent each class label. The class description features are generated once as described in Section 2. The exact prompts used to generate image descriptions, initial predictions, and class description are detailed in Appendix A.
@@ 3.1 Results
We evaluated our method on the following datasets: ImageNet (Deng et al. (2009)), Pets (Parkhi et al. (2012)), Places365 (Zhou et al. (2017)), Food-101 (Bossard et al. (2014)), SUN397 (Xiao et al. (2010; 2016)), Stanford Cars (Krause et al. (2013)), Describable Textures Dataset (DTD) (Cimpoi et al. (2014)), Caltech-101 (Fei-Fei et al. (2004)), CIFAR-10 (Krizhevsky et al. (2009)), and CIFAR-100 (Krizhevsky et al. (2009)). We compared our results against the following zero-shot classification methods: CLIP (Radford et al. (2021)), SLIP (Mu et al. (2022)), PyramidCLIP (Gao et al. (2022)), nCLIP (Zhou et al. (2023)), NLIP (Huang et al. (2023)), UniCLIP (Lee et al. (2022)), ALIP (Yang et al. (2023)), CALIP (Guo et al. (2023)), and CuPL (Pratt et al. (2023)). For CuPL (Pratt et al. (2023)), we utilized Gemini Pro (Gemini Team Google (2023)) for computing class descriptions instead of GPT-3 (Brown et al. (2020)), ensuring a fair comparison with our method, which also employs Gemini Pro. Additionally, it is worth mentioning that the results reported in the original paper of CuPL (Pratt et al. (2023)) use different class labels for the ImageNet dataset (Deng et al. (2009)) compared to the original class labels of the dataset. Consequently, we decided to re-compute the results for CuPL using Gemini Pro with the standard ImageNet class labels.
@@
CLIP (ViT-L/14) (Radford et al. (2021)) | 65.1 | 87.6 | 54.3 | 86.8 | 61.2 | 65.1 | 46.7 | 83.4 | 87.9 | 37.3 | ||||||||||||||||||||
CLIP (ViT-B/32) (Radford et al. (2021)) | 47.0 | 73.4 | 41.4 | 70.8 | 56.5 | 43.8 | 37.0 | 84.4 | 72.9 | 35.6 | ||||||||||||||||||||
CLIP (ViT-B/16) (Radford et al. (2021)) | 54.9 | 66.5 | 37.6 | 80.1 | 58.5 | 52.6 | 37.9 | 85.3 | 80.4 | 36.3 | ||||||||||||||||||||
CLIP (RN50) (Radford et al. (2021)) | 43.0 | 42.7 | 16.0 | 57.4 | 46.3 | 34.3 | 30.5 | 79.1 | 64.7 | 29.7 | ||||||||||||||||||||
CLIP (RN101) (Radford et al. (2021)) | 44.0 | 49.0 | 22.4 | 64.3 | 50.2 | 44.5 | 34.0 | 83.2 | 66.3 | 30.9 | ||||||||||||||||||||
SLIP (Mu et al. (2022)) | 47.9 | 87.5 | 54.2 | 69.2 | 56.0 | 9.0 | 29.9 | 80.9 | 41.6 | - | ||||||||||||||||||||
PyramidCLIP (Gao et al. (2022)) | 47.8 | 81.5 | 53.7 | 67.8 | 65.8 | 65.0 | 47.2 | 81.7 | 83.7 | - | ||||||||||||||||||||
nCLIP (Zhou et al. (2023)) | 48.8 | 83.4 | 54.5 | 65.8 | 59.9 | 18.0 | 57.1 | 73.9 | 33.2 | - | ||||||||||||||||||||
NLIP (Huang et al. (2023)) | 47.4 | 81.9 | 47.5 | 59.2 | 58.7 | 7.8 | 32.9 | 79.5 | 39.2 | - | ||||||||||||||||||||
UniCLIP (Lee et al. (2022)) | 54.2 | 87.8 | 56.5 | 64.6 | 61.1 | 19.5 | 36.6 | 84.0 | 69.2 | - | ||||||||||||||||||||
ALIP (Yang et al. (2023)) | 40.3 | 83.8 | 51.9 | 45.4 | 47.8 | 3.4 | 23.2 | 74.1 | 30.7 | - | ||||||||||||||||||||
CALIP (ViT-B/32) (Guo et al. (2023)) | 60.6 | 76.5 | 44.2 | 77.4 | 58.6 | 56.3 | 42.4 | 87.7 | 86.2 | 36.9 | ||||||||||||||||||||
CALIP (ViT-B/16) (Guo et al. (2023)) | 57.5 | 70.2 | 41.3 | 80.7 | 60.6 | 50.1 | 39.5 | 86.2 | 79.1 | 38.4 | ||||||||||||||||||||
CALIP (RN101) (Guo et al. (2023)) | 50.0 | 49.7 | 24.5 | 67.3 | 51.8 | 39.4 | 35.4 | 84.3 | 70.5 | 33.8 | ||||||||||||||||||||
CuPL (Pratt et al. (2023)) | 66.6 | 86.6 | 57.7 | 89.0 | 65.3 | 63.9 | 49.1 | 90.5 | 80.0 | 39.7 | ||||||||||||||||||||
Tip-Adapter () (Zhang et al. (2022)) | 62.0 | - | - | - | - | - | - | - | - | - | ||||||||||||||||||||
SuS-X () (Udandarao et al. (2023)) | 61.9 | - | - | - | - | - | 50.6 | - | 77.6 | - | ||||||||||||||||||||
Tip-Adapter-F () (Zhang et al. (2022)) | 65.5 | - | - | - | - | - | - | - | - | - | ||||||||||||||||||||
CLIP-Adapter () (Gao et al. (2024)) | 61.3 | - | - | - | - | - | 66.1 | 93.4 | - | - | ||||||||||||||||||||
APE-T () (Zhu et al. (2023)) | 66.1 | - | - | - | - | - | - | - | - | - | ||||||||||||||||||||
@@
Ours (combined) | 73.4 | 93.4 | 70.2 | 93.0 | 70.6 | 76.6 | 58.0 | 89.4 | 90.9 | 43.4 |
For CLIP (Radford et al. (2021)) and CALIP (Guo et al. (2023)), we employed the template “A photo of {class_label}” to encode textual class labels across all datasets, except for Pets (Parkhi et al. (2012)), DTD (Cimpoi et al. (2014)), and Cars (Krause et al. (2013)) datasets. For these exceptions, we employed specific templates: “A photo of {class_label}, a type of pets”, “A photo of {class_label}, a textural category”, and “A photo of {class_label}, a car model”, respectively. This approach was found to enhance results, consistent with prior findings (Radford et al. (2021); Li et al. (2023); Allingham et al. (2023); Popp et al. (2024)).
For consistency, we resize all images to 224224 before processing them with our method, CLIP, CuPL, and CALIP methods. Additionally, we report results from training-free and few-shot learning methods for a comprehensive comparison, including: Tip-Adapter (Zhang et al. (2022)), SuS-X (Udandarao et al. (2023)), CLIP-Adapter (Gao et al. (2024)), and APE-T (Zhu et al. (2023)).
@@
The top-1 accuracy results are reported in Table 1 (see Appendix B for top-5 accuracy results). As can be seen, our method consistently achieves state-of-the-art results when compared with prior work across all datasets in zero-shot image classification methods, and in the majority of datasets when considering other methods (i.e., training-free and few-shot methods). Using the combined class features yields the most promising results across the majority of datasets, with the exception of Caltech-101 (Fei-Fei et al. (2004)), where the best results were achieved using the class description features.
@@
Figure 5 shows visual examples, where we highlight parts of input modalities (image, initial prediction, and image description) based on their contribution to the final predicted class. To emphasize the significance of each input component in the final prediction, we employ a straightforward approach. Specifically, we utilize a 2D sliding kernel that traverses the image, masking out patches of the image. Subsequently, we measure the difference between the initial prediction and the prediction after masking to highlight areas of the image that contribute most significantly to the final prediction. Similarly, we apply this approach to the textual inputs. Given two distinct input texts – namely, the initial prediction and the image description generated by the LLM (Gemini Team Google (2023)) – we utilize a sliding kernel with a stride of one word. We mask out words that match the kernel and quantify their importance in our final prediction. As shown in Figure 5, the three inputs collectively contribute to predicting the final class label. In some cases, one or two inputs exhibit a higher level of influence than the others, as demonstrated in the first, second, and third examples.
@@
It is worth mentioning that, while Gemini’s initial predictions do not match the ground-truth class in the first and second examples, the predictions are contextually sensible. In the first example, Gemini’s prediction was ‘tunnel’ which, while not directly matching any class label in the Places dataset (Zhou et al. (2017)), conceptually aligns with the displayed ‘catacomb’ image as an underground passage. Similarly, in the second example, Gemini’s initial prediction was ‘Snoopy’, which corresponds to the character drawn on the pen shown in the input image. However, ‘Snoopy’ is not one of the ImageNet (Deng et al. (2009)) class labels and the correct class of the shown image in second row of Figure 5 is ‘ballpoint pen’. This behavior of LLMs is the reason we cannot use them directly as image classifiers, because they sometimes do not restrict the output class to the provided list of target classes. However, such behavior might be beneficial to other classification tasks that are not restricted to a specific set of classes. Additional examples are provided in the supplemental materials (Appendix B).

### v3

@@ Abstract
Large language models (LLMs) have been effectively used for many computer vision tasks, including image classification. In this paper, we present a simple yet effective approach for zero-shot image classification using multimodal LLMs. Using multimodal LLMs, we generate comprehensive textual representations from input images. These textual representations are then utilized to generate fixed-dimensional features in a cross-modal embedding space. Subsequently, these features are fused together to perform zero-shot classification using a linear classifier. Our method does not require prompt engineering for each dataset; instead, we use a single, straightforward set of prompts across all datasets. We evaluated our method on several datasets and our results demonstrate its remarkable effectiveness, surpassing benchmark accuracy on multiple datasets. On average, for ten benchmarks, our method achieved an accuracy gain of 6.2 percentage points, with an increase of 6.8 percentage points on the ImageNet dataset, compared to prior methods re-evaluated with the same setup. Our findings highlight the potential of multimodal LLMs to enhance computer vision tasks such as zero-shot image classification, offering a significant improvement over traditional methods.
1 Introduction and related work
@@
Zero-shot image classification aims to categorize images into classes unseen during training, presenting a significant challenge in computer vision. Recent approaches leverage the power of large language models (LLMs) like GPT-4 [3] to generate prompts for target classes, often in conjunction with vision-language models (VLMs) such as CLIP [21] to embed images and text in a common space. Open-vocabulary models like CLIP [21] and VirTex [6] have shown promise in this area due to their ability to generalize to unseen classes. These models learn to match images with captions from vast amounts of image-text data, allowing for dynamic classification without retraining. Early works like DeViSE [8] pioneered the concept of a joint embedding space for images and text, enabling generalization to unseen classes. Approaches such as CLIP [21], based on contrastive learning, and ALIGN [11], employing a two-stage framework, further refined the alignment of image and text representations. More recent approaches for zero-shot image classification (e.g., [20, 17]) have utilized LLMs to generate prompts (i.e., captions or descriptions) for the target classes to further improve the classification accuracy.
@@
Using only the visual content of images to make predictions can reduce accuracy because images alone often do not contain enough information to match the detailed meanings found in text descriptions. To address this limitation, we propose our method to leverage the capabilities of multimodal LLMs to generate rich textual representations of the input images. Multimodal LLMs, such as GPT-4 [3] and Gemini [9], have demonstrated remarkable abilities in various tasks. They are capable of processing and integrating information from various sources such as text, images, and audio. This allows them to perform tasks that were previously challenging, such as generating detailed image descriptions, answering complex visual questions, and even creating realistic images from text. Inspired by these advancements, we utilize a straightforward set of prompts to generate detailed textual descriptions of the input images (see Figure 1), eliminating the need for complex prompt engineering seen in previous works (e.g., [21, 10]). These textual representations are then fused with visual features to perform zero-shot classification.
While zero-shot image classification using large language models (LLMs) initially appeared promising, our experiments demonstrate insufficient accuracy. Even with careful prompt engineering, LLMs frequently predict classes outside the defined target set. Therefore, we propose a refined approach: encoding LLM predictions as input features for a linear classifier. This classifier constrains the output to the target class set, improving prediction accuracy. Our method offers several key advantages:
Improved accuracy by leveraging rich textual information extracted directly from images, significantly boosting classification accuracy.
Constrained LLM predictions to restrict LLM predictions to the desired target classes.
Simplified prompting using a set of universal prompts, eliminating the need for dataset-specific prompts.
Superior performance in a consistent experimental setup, outperforming existing methods across various benchmark datasets. Specifically, our method achieves an average accuracy gain of 6.2 percentage points in ten image classification benchmark datasets and an accuracy increase of 6.8% in the ImageNet dataset [5].
@@
In the following sections, we detail our proposed approach for zero-shot image classification using LLMs (Section 2), present experimental results across ten benchmark datasets and analyze the computational resources used (Section 3), discuss limitations (Section 4), and conclude with remarks and future directions (Section 5).
@@ 2 Method
Given an input image, X, containing object(s) belonging to a single class label from a finite set of class labels, , our objective is to classify X without any dataset-specific training process for image classification. The overview of our method is illustrated in Figure 2. As shown in Figure 2, our approach relies on a cross-modal embedding encoder models (image encoder, , and text encoder, ), trained to learn joint representations of images and text, as demonstrated in prior works, such as [21, 8, 11]. Additionally, we utilize a multimodal LLM, , which is pre-trained on a large corpus of multimodal data. This model, , is designed to generate responses that align with both textual and visual inputs, effectively integrating information from both modalities to make predictions (e.g., Gemini [9]).
@@ 2.1 Class label features
Our zero-shot classifier includes class label features of the target dataset. We first encode the set of class labels, , in some embedding space by the cross-modal text encoder model, , such that refers to the set of encoded class label features, where is a normalized -D class label feature of the class label and is the dimensionality of the embedded features.
@@
Such an encoded feature matrix can be generated using one of three options: (1) directly from the textual class labels, (2) using a human-designed template (e.g., “A photo of {class_label}” [21, 10]), where {class_label} refers to the textual label of one of our classes , or (3) LLM-generated class description(s) [20]. In option (3), the LLM-generated class description(s) are then converted to embedded features, followed by fusion (e.g., averaging) to generate a single embedded feature for each class label in the dataset. Optionally, all features from the three options can be fused together (e.g., averaged) for increased robustness.
@@ 2.2 Cross-modal input features
To predict the final class, we first encode the input image by the cross-modal image encoder model, , to generate the image feature as follows:
where performs vector normalization. Traditionally, the image feature serves as the sole input to prior zero-shot image classifiers [20, 21, 10]. Our method enhances this input by incorporating the LLM, , which generates additional textual-based inputs for our zero-shot image classifier (see Figure 3). To achieve this, we employ an engineered prompt that instructs the LLM to describe the input image, X, and perform initial image classification using the textual names of the class label set . Denoting and as our prompts for image description and initial image classification, respectively, we generate two additional embedded features alongside as follows:
where and refer to the image description feature and initial class prediction feature, respectively. Notably, in our method, unlike prior methods (e.g., [20, 21, 10]), such prompts and do not require dataset-specific engineering for each dataset. Instead, we employ fixed prompts: the first prompt instructs the LLM to provide a generic image description, while the second prompt includes the textual class labels of the target dataset.
| Image classification |
| You are given an image and a list of class labels. Classify the image given the class labels. Answer using a single word if possible. Here are the class labels: {classes} |
| Image description |
| What do you see? Describe any object precisely, including its type or class. |
| Class description |
| 1. Describe what a {class_label} looks like in one or two sentences. |
| 2. How can you identify a {class_label} in one or two sentences? |
| 3. What does a {class_label} look like? Respond with one or two sentences. |
| 4. Describe an image from the internet of a {class_label}. Respond with one or two sentences. |
| 5. A short caption of an image of a {class_label}: |
After generating the three input features (image feature, description feature, and initial prediction feature), we fuse them to generate our final query feature, . One can interpret this fusion as an ensemble of different candidate features to generate a more precise query feature (see Figure 4). We adopted a simple averaging fusion, where the final query feature, , is generated by:
@@ 2.3 Class label prediction
where represents the vector transpose operation to transform into a row vector of shape , and contains the similarity scores of the generated query feature to the class label features in the target dataset. The index of the final predicted class is then computed as that corresponds to the maximum similarity score. The inference process of our method is concisely described in Algorithm 1.
@@ 3 Experiments
In our experiments, we employed Gemini [9] as our multimodal LLM, , for generating image descriptions and initial predictions. We utilized CLIP (ViT-L/14) [21] as our cross-modal embedding encoder models, and , to encode input testing images, image descriptions, and initial predictions generated by Gemini.
We explored four different versions of class label features, M (Equation 1). Specifically, we used CLIP to encode the following representations of class labels: 1) class label names, 2) the text template ‘‘A photo of {class_label}’’, where {class_label} denotes each class label in each dataset, and 3) class descriptions, similar to those generated by CuPL [20], and 4) a combination of the aforementioned three options—akin to the fusion of our input features, we combined the three encoded features of each class and computed their average feature. The class descriptions were produced by prompting Gemini to describe each class label in the dataset 50 times, resulting in 50 different descriptions for each class. Subsequently, we utilized CLIP to encode the 50 class descriptions for each class and compute the average encoded feature vector to represent each class label. The class description features are generated once as described in Section 2.
Table 1 shows the prompts used for each step in our method employing a multimodal LLM (Gemini [9]). The table shows the prompts used to: (1) perform zero-shot image classification with Gemini [9], (2) describe a given testing image, and (3) generate class label descriptions. Both image classification and description can be combined in one prompt. The class descriptions were generated using five prompts, with 10 responses generated for each prompt and class, resulting in 50 class descriptions per class label. To encourage diversity in Gemini’s responses, we set the temperature parameter to a high value of 0.99, as done in [20].
@@ 3.1 Results
We evaluated our method on the following datasets: ImageNet [5], Pets [18], Places365 [27], Food-101 [2], SUN397 [25, 26], Stanford Cars [12], Describable Textures Dataset (DTD) [4], Caltech-101 [7], CIFAR-10 [13], and CIFAR-100 [13]. We compared our results against two baselines using the same backbone as our method (ViT-L/14): (1) CLIP [21] and (2) CuPL [20]. The rationale for selecting these baselines is as follows. First, CLIP is a core component of our method, and comparing against it highlights the impact of our additional features and design choices. Second, CuPL extends CLIP by LLM to enrich the feature representations of target class labels. Since our method also employs an LLM for zero-shot classification, CuPL serves as a strong comparison baseline.
For CLIP [21], we used a fixed text template, identical to our method, to ensure a fair comparison. Specifically, we employed the template “A photo of {class_label}” to encode textual class labels across all datasets, except for Pets [18], DTD [4], and Cars [12] datasets. For these exceptions, we employed specific templates: “A photo of {class_label}, a type of pets”, “A photo of {class_label}, a textural category”, and “A photo of {class_label}, a car model”, respectively. This approach was found to enhance results, consistent with prior findings [21, 14, 1, 19]. For CuPL [20], we replaced GPT-3 [3] with Gemini [9] to generate class descriptions, ensuring a fair comparison with our method, which also utilizes Gemini [9].
Additionally, it is worth noting that the results reported in the original CuPL paper [20] were computed using modified class labels for the ImageNet dataset [5], differing from the dataset’s original labels. To ensure consistency and fairness, we re-evaluated the baseline methods using the standard ImageNet class labels while following the experimental setup described above. Additionally, for uniformity, we resized all images to 224224 before processing them with our method, CLIP [21], and CuPL [20].
@@
| CLIP (ViT-L/14) [21] | 65.1 | 87.6 | 54.3 | 86.8 | 61.2 | 65.1 | 46.7 | 83.4 | 87.9 | 37.3 | ||||||||||||||||||||
| CuPL [20] | 66.6 | 86.6 | 57.7 | 89.0 | 65.3 | 63.9 | 49.1 | 90.5 | 80.0 | 39.7 | ||||||||||||||||||||
@@
| Ours (combined) | 73.4 | 93.4 | 70.2 | 93.0 | 70.6 | 76.6 | 58.0 | 89.4 | 90.9 | 43.4 | ||||||||||||||||||||
@@
The top-1 accuracy results are reported in Table 2 (see supplemental materials for top-5 accuracy results). As can be seen, our method consistently achieves the best accuracy when compared with baseline methods across all datasets. Using the combined class features yields the most promising results across the majority of datasets, with the exception of Caltech-101 [7], where the best results were achieved using the class description features.
@@
Figure 5 shows visual examples, where we highlight parts of input modalities (image, initial prediction, and image description) based on their contribution to the final predicted class. To emphasize the significance of each input component in the final prediction, we employ a straightforward approach. Specifically, we utilize a 2D sliding kernel that traverses the image, masking out patches of the image. Subsequently, we measure the difference between the initial prediction and the prediction after masking to highlight areas of the image that contribute most significantly to the final prediction. Similarly, we apply this approach to the textual inputs. Given two distinct input texts – namely, the initial prediction and the image description generated by the LLM (Gemini [9]) – we utilize a sliding kernel with a stride of one word. We mask out words that match the kernel and quantify their importance in our final prediction. As shown in Figure 5, the three inputs collectively contribute to predicting the final class label. In some cases, one or two inputs exhibit a higher level of influence than the others, as demonstrated in the first, second, and third examples.
@@
It is worth mentioning that, while Gemini’s initial predictions do not match the ground-truth class in the first and second examples, the predictions are contextually sensible. In the first example, Gemini’s prediction was ‘tunnel’ which, while not directly matching any class label in the Places dataset [27], conceptually aligns with the displayed ‘catacomb’ image as an underground passage. Similarly, in the second example, Gemini’s initial prediction was ‘Snoopy’, which corresponds to the character drawn on the pen shown in the input image. However, ‘Snoopy’ is not one of the ImageNet [5] class labels and the correct class of the shown image in second row of Figure 5 is ‘ballpoint pen’. This behavior of LLMs is the reason we cannot use them directly as image classifiers, because they sometimes do not restrict the output class to the provided list of target classes. However, such behavior might be beneficial to other classification tasks that are not restricted to a specific set of classes. Additional examples are provided in the supplemental materials.